# Fully Spiking Neural Network for Legged Robots

### v2

@@ IV Experiments
The objectives of our experiments are the followings: i) To validate the feasibility of SNNs on robots operating in environments with high latitudes and intricate dynamics models. ii) Verify the benefits of SNNs over ANNs in terms of ultra-high flatness control. iii) Identify rewards that SNNs excel on rl tasks and make improvements on poorly performing rewards. We assessed our approach using the Isaac Gym platform, which offers precise physics simulation and a versatile, customizable framework for robot models and environments.
We primarily evaluated the performance of the following robots: A1, Cassie [41], and MIT Humanoid [42] where each SNN is trained within $1,500,000$ iterations until convergence. The ultimate assessment involves utilizing visual representations and metrics, including base linear velocity in the x and y axes, angular velocity in the yaw axis, mean episode length, and position torque are used for evaluation.
IV-A Simulation Setup
TABLE I: Ranges of the environmental parameters
Parameters Training range Testing range
Friction [0.005, 4.5] [0.004, 6.0]
$K_{p}$ [50, 60] [45, 65]
$K_{d}$ [0.4, 0.8] [0.3, 0.9]
Payload(Kg) [0, 6] [0, 7]
Center of Mass(cm) [-0.15, 0.15] [-0.18, 0.18]
Motor Strength [0.90, 1.10] [0.88, 1.22]
Re-sample Probability 0.004 0.01
@@ IV-A Simulation Setup
In order to establish a varied array of environments, we incorporated multiple URDF files from recent studies. The files contain various models, including A1, Anymal-b, Anymal-c, Cassie, MIT Humanoid, and others. Once imported, we utilize the functionalities of the Isaac Gym simulator to generate novel environments for each of these models. To introduce diversity, we make use of the built-in fractal terrain generator offered by the simulator. This terrain generator facilitates the creation of diverse types of terrain, encompassing flat terrains as well as rugged terrains with a variety of topographical characteristics. The policy functions at a control frequency of up to 500Hz, attributed to our SNN-based approach. This capability enables the attainment of high-frequency control, leading to prompt and precise adjustments to the system. In Table I listed a comprehensive list of environmental variations and their respective ranges is presented.
IV-B Performances of High Frequency Control using SNNs
@@
We conducted tests on the robots mentioned above specifically for linear and angular velocity tracking tasks. In these tests, we considered tracking velocity (linear and angular) as the primary positive rewards. Additionally, we incorporated penalties for base height that is too low, excessive acceleration, and instances where the robot falls, etc. For A1, we conducted training and testing in several terrain environments, including pyramid stairs like terraces (upstairs and downstairs), pyramids with sloping surfaces, hilly terrains, terrains with discrete obstacles, and terrains covered with stepping stones. On the other hand, Cassie is solely trained in a trapezoidal pyramid environment and MIT Humanoid in a plain terrain.
IV-B1 A1
To explore the benefits of SNNs in high-frequency control scenarios, we intentionally increased the simulation environment’s time step (dt) to 2.5 times the default ANNs task, achieving a frequency of 500Hz. Our goal is to determine if SNN outperforms ANN in real-world environments that demand high-frequency control by comparing their performance.
Figure 5: The four graphs depicting the robot’s x-axis linear velocity, y-axis linear velocity, yaw-axis angular velocity, and DOF position vividly demonstrate the successful accomplishment of the task through the implementation of our method.
The investigation is motivated by the limitations robots face when using ANN control algorithms due to energy constraints. As a result, these robots can typically only achieve a policy inference frequency of 100Hz, significantly lagging behind the motors’ execution frequency. Instead, SNN has the potential to improve policy inference quality and provide real-time control performance by leveraging its energy-efficient nature and deployment on specialized processors. If SNN can match or outperform ANN in high-frequency control, it would demonstrate SNN’s superiority in real-world environments.
In Figure 5, the effectiveness of our method under high-frequency control is clearly demonstrated. The A1 robot demonstrates impressive performance in tracking velocity x, accurately following the desired trajectory. Considering the impact of complex terrain conditions, all velocities varied within a margin of $17\%$ around the designated value. Although occasional spikes occur, indicating moments when the robot deviates from the existing policy framework, these instances can be attributed to the robot’s inclination to explore alternative approaches when faced with challenging terrain. This emphasizes the robustness of our method in tackling challenging terrains.
IV-B2 Cassie
@@
Figure 6: The first image showcases Cassie’s remarkable capability to conquer complex terrain, as indicated by the terrain level value nearing 6. Additionally, the second figure demonstrates Cassie’s impeccable tracking of the angular velocity of the yaw axis, exemplifying its capacity to maintain body stability while traversing complex terrain.
In Cassie’s experiments, a trapezoidal pyramid environment was utilized, with terrain level indicates the absolute value of the robot’s elevation displacement (a value of 6 signifies reaching the top of the sixth-order pyramid). The training results of the Cassie robot have proven to be highly successful, as shown in Figure 6. The stability of the angular velocity along the yaw axis is notable. This stability is crucial for maintaining effective control and balance while navigating diverse terrains. Furthermore, the robot’s impressive ability to reach the highest terrain levels demonstrates its adaptability in conquering rugged landscapes. These findings highlight the Cassie robot’s ability to navigate difficult terrains effectively using SNNs at a high control frequency while maintaining stability. Consequently, the feasibility of our approach is confirmed, supporting its potential deployment in complex and dynamic real-world scenarios.
IV-B3 MIT Humanoid
@@
The training of MIT Humanoid also proved to be highly successful, showcasing the effectiveness of our spike-based approach. While it did take slightly longer to train compared to the traditional ANNs, the results obtained are equally impressive. In fact, they even surpassed the ANN in certain individual metrics, as clearly depicted in Figure 7. These findings strongly suggest that the SNN possesses inherent advantages when it comes to control robustness. Furthermore, it enables agents to not only thrive but also endure in their environment for an extended duration. This highlights the potential of SNN as a powerful tool in enhancing agent performance and longevity.
The performance demonstrated by our approach, whether through A1 and Cassie’s agile traversal of challenging terrains or the MIT Humanoid’s unrestricted running, is undeniably superior.
IV-C Continuity Comparison
Compared to ANN, SNN also shows stronger robustness. This is because in SNNs the signaling relies on thresholding, leading to the direct exclusion of some noise. The stochastic behavior of spiking neuron dynamics also enhances the robustness of the network to external noise. We added Gaussian noise to the robot’s movement commands, and tested the performance of the networks with two strategies, ANN and SNN, for sigma of 0.1, 0.2, and 0.3, respectively, using the following of the robot’s y-axis linear velocity as an index. The experiment (Figure 8) proves that as the noise increases, the SNN’s ability to resist noise becomes more and more obvious, and its measurements become closer to the desired value and are accompanied by less fluctuations. On the other hand, the situation of ANN is not optimistic, at sigma of 0.3, the trained strategy is completely insufficient to support the robot to walk normally, and falls occurred. Under the same noise interference, the robot with SNN as the policy network can still walk smoothly.
@@ IV-C Continuity Comparison
IV-D Estimation of Energy Consumption
As mentioned earlier, one of the primary advantages of our SNN policy is its minimal energy usage. The assessment of energy consumption in a SNN is complex because the floating-point operations (FLOPs) in the initial encoder layer are MAC, whereas all other Conv or FC layers are AC. Building upon prior research[43, 44, 45, 46] conducted by SNN, it is assumed that the data utilized for various operations is represented in 32-bit floating-point format in 45nm technology[45], with $E_{MAC}$ = 4.6pJ and $E_{AC}$ = 0.9pJ. The energy consumption equations simulated for SNN are provided as follows:
@@ IV-D Estimation of Energy Consumption
Experimental results (Table II) demonstrate that our approach offers a significant energy efficiency improvement compared to the conventional ANN architecture. Specifically, it achieves energy savings of $95.65\%$, $81.49\%$, and $60.34\%$ at T = 1, 2 and 3, respectively.
TABLE II: Energy Comparison(×$\times 10^{-6}$ mJ)
Method Actor(T=1) Actor(T=2) Actor(T=3)
ANN Model 85.96 85.96 85.96
SNN Model(ours) $3.35$ $15.36$ $35.22$
Energy Saving 96.10% 82.13% 59.03%
V Conclusion
This study presents the integration of SNN with history trajectory and imitation learning, which achieves performance comparable to ANNs, highlights the versatility of SNNs in policy gradient-based DRL algorithms. This opens up new horizons for application of SNNs in various reinforcement learning tasks, including continuous control.
Additionally, our approach offers significant advantages in energy efficiency, addressing signal noise, and high-frequency control. These advantages are significant for improving structural integrity and robustness in practical situations. It also has potential to reduce costs related to robot development and enables the implementation of advanced sensing systems in robotic platforms.

### v3

@@ III Experiments
The objectives of our experiments are the followings: i) To validate the feasibility of SNNs on robots operating in environments with high dimensions and intricate dynamics models. ii) Verify the benefits of SNNs over ANNs in terms of ultra-high frequency control. We assessed our approach using the Isaac Gym platform, which is a simulation platform specifically created for robotics applications. We primarily evaluated the performance of the following robots: A1 [35], Cassie [35], and MIT Humanoid [36] where each SNN is trained within $1,500,000$ iterations until convergence.
Figure 5: Four graphs illustrate the exceptional performance of the robot in command-following task.
III-A Simulation Setup
@@
In order to establish a varied array of environments, we incorporated multiple URDF files from recent studies. The files contain various models, including A1, Anymal-b, Anymal-c, Cassie, MIT Humanoid, and others. Once imported, we utilize the built-in fractal terrain generator of the Isaac Gym simulator to generate various environments for each of these models to introduce diversity. The policy functions at a control frequency of 500 Hz due to our SNN-based method, enabling quick and accurate system adjustments.
III-B Performances of High Frequency Control using SNNs
@@
We tested the aforementioned robots specifically on linear and angular velocity tracking tasks, using tracking velocity (linear and angular) as the primary reward. Penalties were applied for low base height, excessive acceleration, and instances of the robot falling, etc. For A1, we conducted training and testing in several terrain environments, including pyramid stairs like terraces (upstairs and downstairs), pyramids with sloping surfaces, hilly terrains, terrains with discrete obstacles, and terrains covered with stepping stones. On the other hand, Cassie is solely trained in a trapezoidal pyramid environment and MIT Humanoid in a plain terrain.
III-B1 A1
To explore the benefits of SNNs in high-frequency control, we increased the simulation environment’s time step (dt) to 2.5 times that of the default ANNs task, achieving 500 Hz. ANNs are energy-constrained and typically reach only 100 Hz, lagging behind motor execution frequency. In contrast, SNNs may improve policy inference quality and enable real-time control due to their energy efficiency. If SNNs can match or exceed ANNs in high-frequency control, it would demonstrate their superiority in real-world environments. Figure 5 shows the A1 robot effectively tracking velocity x and following the desired trajectory in complex terrain, with all velocities varying within $17\%$ of the designated value.
III-B2 Cassie
@@
Figure 6: The first image showcases Cassie’s remarkable capability to conquer complex terrain, as indicated by the terrain level value nearing 6. Additionally, the second figure demonstrates Cassie’s impeccable tracking of the yaw axis’s angular velocity, highlighting its stability while traversing complex terrain.
@@
In Cassie’s experiments, the terrain level indicates the absolute value of the robot’s elevation displacement (a value of 6 signifies reaching the top of the sixth-order pyramid). Figure 6 highlights the stability of angular velocity, essential for maintaining effective control and balance while navigating diverse terrains. The robot’s capability to reach the highest terrain levels further demonstrates its adaptability in conquering rugged landscapes.
III-B3 MIT Humanoid
The MIT Humanoid training showcased the effectiveness of our spike-based approach. While it took longer to train than traditional ANNs, the results were impressive. In fact, they even surpassed the ANN in certain individual metrics, as clearly depicted in Figure 7. These findings strongly suggest that the SNN possesses advantages when it comes to control robustness. The performance demonstrated by our approach, whether through A1 and Cassie’s agile traversal of challenging terrains or the MIT Humanoid’s unrestricted running, is undeniably superior.
III-C Ablation Experiment
Figure 8: Results indicate that SNN-based DRL achieved performance comparable to ANN in high-frequency control scenarios, and our method yields the highest rewards. Plots are smoothed for clarity.
We chose the Proximal Policy Optimization (PPO) algorithm [37] to evaluate our approach on three different robotic structures, assessing its universality. Figure 8 shows that while the SNN-based DRL algorithm converges more slowly than the ANN-based DRL algorithm, it ultimately achieves comparable training results.
III-D Continuity Comparison
SNNs exhibit greater robustness than ANNs because their thresholding mechanism filters noise, and their stochastic dynamics improve resilience to external disturbances. We added Gaussian noise to the robot’s movement commands, and tested the performance of the networks with two strategies, ANN and SNN, for sigma of 0.1, 0.2, and 0.3, respectively, using the following of the robot’s y-axis linear velocity as an index. The experiment (Figure 8) shows that as noise increases, SNN demonstrates greater resilience, with measurements aligning closely to desired values and exhibiting fewer fluctuations. In contrast, ANN performs poorly; at a sigma of 0.3, it fails to support normal robot movement, leading to falls. The robot using the SNN, however, walks smoothly under the same noise conditions.
@@ III-D Continuity Comparison
III-E Estimation of Energy Consumption
A key advantage of our SNN policy is its minimal energy usage. The assessment of energy consumption in a SNN is complex because the floating-point operations (FLOPs) in the initial encoder layer are MAC, whereas all other Conv or FC layers are AC. Building upon prior research[38, 39, 40, 41] conducted by SNN, it is assumed that the data utilized for operations is represented in 32-bit floating-point format in 45nm technology[40], with $E_{MAC}=$ 4.6pJ and $E_{AC}=$ 0.9pJ. The energy consumption equations for SNN are shown below:
@@ III-E Estimation of Energy Consumption
Experimental results (Table I) indicate a significant energy efficiency improvement over conventional ANN architecture. Specifically, it achieves energy savings of $96.01\%$, $81.99\%$, and $58.86\%$ at $T_{i}$ = 1, 2 and 3, respectively. Each $T_{i}$ in the table represents the final stage, with $i$ set to $3$. The time dimension decreases by $1$ after each stage. The temporal shrinking between stages mainly involves simple fully connected layers, which are negligible compared to the final classifier. As the auxiliary classifier is not used during inference, overall energy consumption approximates that of the SNN main classifier.
TABLE I: Energy Comparison(×$\times 10^{-6}$ mJ)
Method Actor($T_{i}$=1) Actor($T_{i}$=2) Actor($T_{i}$=3)
ANN Model 86.27 86.27 86.27
SNN Model(ours) $3.44$ $15.54$ $35.49$
Energy Saving 96.01% 81.99% 58.86%
IV Conclusion
This study presents the integration of a lightweight population coded SNN trained in a multi-stage method with history trajectory and imitation learning, which achieves performance comparable to ANNs, highlighting the versatility of SNNs in policy gradient-based DRL algorithms. This opens up new horizons for application of SNNs in various reinforcement learning tasks, including continuous, high dimensional control. Additionally, our approach offers significant advantages in energy efficiency, addressing signal noise, and high-frequency control. These advantages are significant for improving structural integrity and robustness in practical situations. It also has potential to reduce costs related to robot development and enables the implementation of advanced sensing systems in robotic platforms.